OpenStudy (anonymous):

how to factor 6cos(x)sin(x) - 2cos(x)=0

OpenStudy (anonymous):

factor out cos x

OpenStudy (anonymous):

or better yet, 2 cosx. It will be (2cosx)(3sinx-1)=0

OpenStudy (anonymous):

Yes, that's what I got. Now, how do you set each factor to zero??

OpenStudy (anonymous):

2 cos(x)=0 and 3sin(x)-1=0 ?

OpenStudy (anonymous):

@tfarand for the first part. 2cosx=0 divide both sides by 2, so cosx=0/2 but 0/2 is just 0 so x=cos^-1(0)= 90 degrees (pi/) or 270 degrees (3pi/2) On the other hand, 3sinx-1=0 transpose -1. it will be 3sinx=1 divide both sides by 3 sinx=1/3 x=sin^-1(1/3) x= ?

OpenStudy (anonymous):

That's exactly what I got, but there aren't any solutions with sinx=1/3

OpenStudy (anonymous):

i guess you must use your calculator. :)

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (anonymous):

:)